# Implicit Discriminative Knowledge Learning for Visible-Infrared Person Re-Identification

### v2

@@ 2 Related Work
Visible Modality Person ReID. Person ReID has received increasing success in recent years, which aims to implement pedestrian retrieval between visible images. It has suffered huge challenges including various viewpoints, illuminations, postures and so on. To solve these problems, most methodologies [25, 24, 48, 39, 20] are designed to obtain unified intra-class and discriminative inter-class representation by training CNN network. To further enhance the distinguishability of features, Sun et al. [27] proposed aligning part feature directly instead of using external cues. He et al. [9] caught the hot spot of transformer and first proposed the transformer variant applying on single-modality person ReID. Simultaneously, person ReID has developed some significant branches yet, such as Unsupervised Domain Adaptive Person ReID (UDA-ReID) and Domain Generalization Person ReID (DG-ReID) [41]. Although existing methods have made wide progress in visible modality person ReID, they suffer performance degradation when applied to visible-infrared person ReID due to the severe cross-modality discrepancy.
@@ 3 Methodology
where $E_{sp}$ and $E_{sh}$ denote the specific and the shared feature extractor by employing ResNet-50, respectively. Among the sturctures of ResNet-50, the global average pooling replaced by Gem pooling[24] which is a pooling operation between maximum pooling and average pooling. And $\Theta$ is the shallow layer parameters with the first three blocks of ResNet-50, $\Psi$ and $\Phi$ is the deep layer parameters of different branches with the last two blocks of ResNet-50.
@@ 3.1 Modality Confuser and Discriminator
$\mathcal{L}_{C}=\sum_{j=1}^{K}\mathcal{L}_{Cj},\quad\mathcal{L}_{D}=\sum_{j=1}^{K}\mathcal{L}_{Dj}$ (4)
@@
$\mathcal{L}_{b}=\mathcal{L}_{sh}+\mathcal{L}_{sp}$ (6)
@@ 3.2 Information Purifier
$\mathcal{L}_{e}=\operatorname{Softplus}\left(h\left(C_{sp}\left(\boldsymbol{f}^{d+}_{sp}\right)\right)-h(C_{sp}(\boldsymbol{f}_{sp}))\right)$ (10)
@@
Finally, an alternating integration strategy is employed to extract the distinctive information from $\widehat{\mathbf{F}}_{sp}^{m-}$ by applying $\mathbf{m}_{e}$ and the invariant information from $\mathbf{F}_{sp}^{d+}$ by applying $\mathbf{m}_{r}$. Thus, The two sets of features extracted in this way both demonstrate smaller modality differences and stronger implicit discriminative information. By integrating them, we derive the output purified modality-specific feature $\widetilde{\boldsymbol{F}}_{sp}$ as:
$\widetilde{\boldsymbol{F}}_{sp}=\mathbf{m}_{e}\odot\widehat{\boldsymbol{F}}_{sp}^{m-}+\mathbf{m}_{r}\odot\boldsymbol{F}_{sp}^{d+},$ (12)
@@
$\mathcal{L}_{ip}=\mathcal{L}_{e}+\mathcal{L}_{r}+\mathcal{L}_{reid}\left(\tilde{\boldsymbol{f}}_{sp}\right)$ (13)
@@ 3.3.1 Triplet Graph Structure Alignment (TGSA)
$\boldsymbol{A}_{(a;b)}$ =$=\left\{\alpha_{ij}\right\}_{i,j\in N}$ $=\frac{\exp\left(D\left(\boldsymbol{f}_{a}^{i},\boldsymbol{f}_{b}^{j}\right)\right)}{\sum_{k\in N}\exp\left(D\left(\boldsymbol{f}_{a}^{i},\boldsymbol{f}_{b}^{k}\right)\right)},$ (15)
@@
Specifically, the triplet graph structure alignment loss $\mathcal{L}_{tgs}$ is developed for cross modality ReID to align two different modality types, enabling them to conform to the same graph structure distribution and reduce modality discrepancy. This loss encompasses two self-modal affinity matrices and one cross-modal affinity matrix, with the latter ensuring the overall consistency of the graph structure’s distribution as shown in Fig. 3. These three matrices are aligned pairwise through the utilization of the Kullback-Leibler (KL) divergence. Therefore, the alignment loss $\mathcal{L}_{tgsa}$ of two distinct modality types is defined as:
@@
In order to convey the discriminative modality-specific knowledge to shared feature on feature-level, the distillation loss across two branches on homogeneous features through TGSA can be formulated as:
@@ 3.3.2 Class Sementic Alignment (CSA)
where $C_{sp}$ and $C_{sh}$ is the modality-specific and modality-shared classifier separately. And the logit $\boldsymbol{Z}_{sp}$ and $\boldsymbol{Z}_{sh}\in\mathbb{R}^{2N\times C}$ of specific and shared branch both contain visible modality $\boldsymbol{Z}_{sp,V}$and infrared modality $\boldsymbol{Z}_{sp,I}\in\mathbb{R}^{N\times C}$, $C$ is the total number of train dataset identities.
@@
$\mathcal{L}_{csa}=\sum_{i=1}^{N}\left(\text{KL}\left(\boldsymbol{Z}_{sh,V}^{i},\boldsymbol{Z}_{sp,V}^{i}\right)+\text{KL}\left(\boldsymbol{Z}_{sh,I}^{i},\boldsymbol{Z}_{sp,I}^{i}\right)\right)$ (19)
@@ 3.4 Modality Discrepancy Reduction (MDR)
$\mathcal{L}_{mdr}=\mathcal{L}_{tgsa}^{(sh,V;sh,I)}+\sum_{i=1}^{N}\text{KL}\left(\boldsymbol{Z}_{sh,V}^{i},\boldsymbol{Z}_{sh,I}^{i}\right),$ (20)
@@ 3.5 Optimization
$\mathcal{L}_{total}=\mathcal{L}_{b}+\lambda_{1}\mathcal{L}_{ip}+\lambda_{2}\mathcal{L}_{tgsa}+\lambda_{3}\mathcal{L}_{csa}+\mathcal{L}_{mdr}\$ (21)
@@ 4.2 Comparison with State-of-the-art Methods
Comparison on SYSU-MM01 and dataset. The comparison experimental results is shown in Tab. 1 which displays the proposed IDKL method outperforms existing cutting-edge methods. Specifically, the IDKL method achieves the accuracy of 81.42% rank-1 and 79.85% map with single-shot all search protocol , while the accuracy of 87.14% rank-1 and 89.37% map with single-shot indoor search protocol. The compared SOTAs include various base methods, i.e., for the methods of learning the shared feature through network and loss function immediately, which contain D-HSME [8], NFS [2] , CMT [14], DEEN[14] and MCLNet [7]. Based on graph structure augment method DDAG [36] and the mutual learning by logits skill (PIC [47] and MPANet [33]). Comparing with several other based modality-specific methods cm-SSFT [40], MUN[21] and MSCLNet[45], our results outperform them by a margin.

### v3

@@ 2 Related Work
Visible Modality Person ReID. Person ReID has received increasing success in recent years, which aims to implement pedestrian retrieval between visible images. It has suffered huge challenges including various viewpoints, illuminations, postures and so on. To solve these problems, most methodologies [25, 24, 48, 39] are designed to obtain unified intra-class and discriminative inter-class representation by training CNN network. To further enhance the distinguishability of features, Sun et al. [27] proposed aligning part feature directly instead of using external cues. He et al. [9] caught the hot spot of transformer and first proposed the transformer variant applying on single-modality person ReID. Simultaneously, person ReID has developed some significant branches yet, such as Unsupervised Domain Adaptive Person ReID (UDA-ReID) and Domain Generalization Person ReID (DG-ReID) [41]. Although existing methods have made wide progress in visible modality person ReID, they suffer performance degradation when applied to visible-infrared person ReID due to the severe cross-modality discrepancy.
@@ 3 Methodology
where $E_{sp}$ and $E_{sh}$ denote the specific and the shared feature extractor by employing ResNet-50, respectively. Among the sturctures of ResNet-50, the global average pooling replaced by Gem pooling[24, 39] which is a pooling operation between maximum pooling and average pooling. And $\Theta$ is the shallow layer parameters with the first three blocks of ResNet-50, $\Psi$ and $\Phi$ is the deep layer parameters of different branches with the last two blocks of ResNet-50.
@@ 3.1 Modality Confuser and Discriminator
$\mathcal{L}_{C}=\sum_{j=1}^{K}\mathcal{L}_{Cj},\quad\mathcal{L}_{D}=\sum_{j=1}^{K}\mathcal{L}_{Dj}.$ (4)
@@
$\mathcal{L}_{b}=\mathcal{L}_{sh}+\mathcal{L}_{sp}.$ (6)
@@ 3.2 Information Purifier
$\mathcal{L}_{e}=\operatorname{Softplus}\left(h\left(C_{sp}\left(\boldsymbol{f}^{d+}_{sp}\right)\right)-h(C_{sp}(\boldsymbol{f}_{sp}))\right),$ (10)
@@
Finally, an alternating integration strategy is employed to extract the distinctive information from $\widehat{\mathbf{F}}_{sp}^{m-}$ by applying $\mathbf{m}_{e}$ and the invariant information from $\mathbf{F}_{sp}^{d+}$ by applying $\mathbf{m}_{r}$. Thus, The two sets of features extracted in this way both demonstrate smaller modality differences and stronger implicit discriminative information. By integrating them, we derive the purified modality-specific feature $\widetilde{\boldsymbol{F}}_{sp}$ as:
$\widetilde{\boldsymbol{F}}_{sp}=\mathbf{m}_{e}\odot\widehat{\boldsymbol{F}}_{sp}^{m-}+\mathbf{m}_{r}\odot\boldsymbol{F}_{sp}^{d+}.$ (12)
@@
$\mathcal{L}_{ip}=\mathcal{L}_{e}+\mathcal{L}_{r}+\mathcal{L}_{reid}\left(\tilde{\boldsymbol{f}}_{sp}\right).$ (13)
@@ 3.3.1 Triplet Graph Structure Alignment (TGSA)
$\boldsymbol{A}_{(a;b)}=\left\{\alpha_{ij}\right\}_{i,j\in N}=\frac{\exp\left(D\left(\boldsymbol{f}_{a}^{i},\boldsymbol{f}_{b}^{j}\right)\right)}{\sum_{k\in N}\exp\left(D\left(\boldsymbol{f}_{a}^{i},\boldsymbol{f}_{b}^{k}\right)\right)},$ (15)
@@
Specifically, the triplet graph structure alignment loss $\mathcal{L}_{tgsa}$ is developed for cross modality ReID to align two different modality types, enabling them to conform to the same graph structure distribution and reduce modality discrepancy. This loss encompasses two self-modal affinity matrices and one cross-modal affinity matrix, with the latter ensuring the overall consistency of the graph structure’s distribution as shown in Fig. 3. These three matrices are aligned pairwise through the utilization of the Kullback-Leibler (KL) divergence. Therefore, the alignment loss $\mathcal{L}_{tgsa}$ of two distinct modality types is defined as:
@@
In order to convey the discriminative implicit modality-specific knowledge to shared feature on feature-level, the distillation loss across two branches on homogeneous features through TGSA can be formulated as:
@@ 3.3.2 Class Sementic Alignment (CSA)
where $C_{sp}$ and $C_{sh}$ is the modality-specific and modality-shared classifier separately. And the logit $\boldsymbol{Z}_{sp}$ and $\boldsymbol{Z}_{sh}\in\mathbb{R}^{2N\times C}$ of specific and shared branch both contain visible modality and infrared modality, $C$ is the total number of train dataset identities.
@@
$\mathcal{L}_{csa}=\sum_{i=1}^{N}\left(\text{KL}\left(\boldsymbol{Z}_{sh,V}^{i},\boldsymbol{Z}_{sp,V}^{i}\right)+\text{KL}\left(\boldsymbol{Z}_{sh,I}^{i},\boldsymbol{Z}_{sp,I}^{i}\right)\right).$ (19)
@@ 3.4 Modality Discrepancy Reduction (MDR)
$\mathcal{L}_{mdr}=\mathcal{L}_{tgsa}^{(sh,V;sh,I)}+\sum_{i=1}^{N}\text{KL}\left(\boldsymbol{Z}_{sh,V}^{i},\boldsymbol{Z}_{sh,I}^{i}\right).$ (20)
@@ 3.5 Optimization
$\mathcal{L}_{total}=\mathcal{L}_{b}+\lambda_{1}\mathcal{L}_{ip}+\lambda_{2}\mathcal{L}_{tgsa}+\lambda_{3}\mathcal{L}_{csa}+\mathcal{L}_{mdr},\$ (21)
@@ 4.2 Comparison with State-of-the-art Methods
Comparison on SYSU-MM01 dataset. The comparison experimental results is shown in Tab. 1 which displays the proposed IDKL method outperforms existing cutting-edge methods. Specifically, the IDKL method achieves the accuracy of 81.42% rank-1 and 79.85% map with single-shot all search protocol , while the accuracy of 87.14% rank-1 and 89.37% map with single-shot indoor search protocol. The compared SOTAs include various base methods, i.e., for the methods of learning the shared feature through network and loss function immediately, which contain D-HSME [8], NFS [2] , CMT [14], DEEN[43] and MCLNet [7]. Based on graph structure augment method DDAG [36] and the mutual learning by logits skill (PIC [47] and MPANet [33]). Comparing with several other based modality-specific methods cm-SSFT[21] , MUN[40] and MSCLNet[45], our results outperform them by a margin.